A particle is thrown vertically upward from ground at 55 m/s. Distance travelled by the body in 6th second is (g = 10 m/s^2)
1) 2.5 m
2) 5 m
3) 10 m
4) 1.25 m

Grade:12th pass

1 Answers

Susmita
5 years ago
It means distance travelled between t=5s to t=6s.
First of all we have to check if the body reaches its maximum height before 6s or not.
At maximum height velocity v=0.
v=u-gt=55-10t
Or,0=55-10t
t=5.5s
So after 5.5 second the body starts falling back.
At t=5s the velocity of the body is
v'=u-gt=55-10×5=5m/s
From t=5 to t=5.5s the body travels upward.The distance covered
h=v't-gt2/2
Here  t=5.5-5=0.5s.
h=5×0.5-10×0.52/2=1.25 m
From t=5.5 to t=6s the body falls downward.at t=5.5 velocity is zero.so distance covered in this 0.5 s is
h=ut+gt2/2=0+10×0.52/2=1.25m
Adding these two distance covered between 5s to 6s is 2.5
If it helps please approve